# Optimal Neural Compressors for the Rate-Distortion-Perception Tradeoff

### v1

@@ 4 Achieving the Fundamental Limits
We now theoretically analyze the performance of SD-LTC and PD-LTC on the Gaussian source from an information-theoretic perspective. Specifically, we describe the RDP tradeoff asymptotically achievable by SD-LTC and PD-LTC. While the operational rate and distortion are given by the per-dimension versions of what was described in Sec. 3, the operational perception used is per-dimension squared 2-Wasserstein distance. We leave full proofs to Appendix. C.
@@ 4.1 Infinite Shared Randomness
The RDP function for Gaussian source $P_{X}=\mathcal{N}(0,\sigma^{2})$, squared-error distortion $\operatorname{\Delta}(x,\hat{x})=(x-\hat{x})^{2}$, and squared 2-Wasserstein distance $\delta(P_{X},P_{\hat{X}})=W_{2}^{2}(P_{X},P_{\hat{X}})$ is given by
@@
Remark 4.3.
The Gaussian RDP function under perfect realism ($P=0$) is given by
$R(D,0)=\frac{1}{2}\log\frac{\sigma^{4}}{D\left(\sigma^{2}-\nicefrac{{D}}{{4}}\right)}.$ (17)
The following theorem shows that $R(D,P)$ is achievable with SD-LTCs, and is an extension of Li et al. (2011), who addressed the $P=0$ case.
Theorem 4.4 (Optimality of SD-LTC for Gaussian sources).
Let $X_{1},X_{2},\dots\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\mathcal{N}(0,\sigma^{2})$. For any $P$ and $D$ satisfying $0\leq P\leq\sigma^{2}$ and $0<D\leq 2\sigma^{2}$, there exists a sequence of SD-LTCs $\{(g_{a}^{(n)},g_{s}^{(n)},\Lambda^{(n)})\}_{n=1}^{\infty}$ such that
@@
$\lim_{n\rightarrow\infty}\frac{1}{n}W_{2}^{2}(X^{n},\hat{X}^{n})\leq P,$ (20)
@@
The proof of Thm. 4.4 relies on a sphere-bound-achieving sequence of lattices with scalar transforms $g_{a}$, $g_{s}$. As dimension grows, the dither $\bm{u}$ becomes Gaussian-like, and the latent dithered LQ acts like a Gaussian channel, imposing a joint Gaussian relationship between $X$ and $\hat{X}$ as desired by the RDP solution, see Remark 4.2.
@@ 4.2 No Shared Randomness
We now consider the case when the encoder and decoder do not have access to any shared randomness. For simplicity and ease of presentation, we consider the regime of near-perfect perception. This is the only regime of perception that has been studied in information theory before. First, let us define the following function:
$R_{p}(D,0):=\frac{1}{2}\log\frac{2\sigma^{4}-D\left(\sigma^{2}-\nicefrac{{D}}{{4}}\right)}{D\left(\sigma^{2}-\nicefrac{{D}}{{4}}\right)}.$ (21)
The following theorem shows that a rate of $R_{p}(D,0)$, distortion $D$, and perception 00 is achievable with PD-LTCs.
Theorem 4.6 (PD-LTC achieves $R_{p}(D,0)$ for Gaussian sources).
@@
$\lim_{n\rightarrow\infty}\frac{1}{n}H(Q_{\Lambda^{(n)}}(g_{a}^{(n)}(X^{n})))=R_{p}(D,0),$ (22)
@@
where $\hat{X}^{n}=g_{s}^{(n)}(Q_{\Lambda^{(n)}}(g_{a}^{(n)}(X^{n}))+\bm{u})$, and $\bm{u}\sim\mathrm{Unif}(\mathcal{V}_{0}(\Lambda^{(n)}))$.
@@
Shared randomness is not available in PD-LTC and the rate $R_{p}(D,0)$ that we prove achievable in Thm. 4.6 is lower bounded by the RDP function $R(D,0)$. Let
$R_{0}(D,0)=\frac{1}{2}\log\frac{2\sigma^{2}}{D}.$ (25)
Figure 3: Comparing Gaussian RDP limits, $\sigma^{2}=1$.
One can verify that $R_{p}(D,0)\approx R_{0}(D,0)$ when $D\ll\sigma^{2}$ and $R_{p}(D,0)\approx R(D,0)$ when $D\approx 2\sigma^{2}$; see Fig. 3. Prior work (Chen et al., 2022) has established that $R(D,0)$ is achievable even without shared randomness for weak-sense perception (see Sec. 2.2). On the other hand, it is proved in Wagner (2022) that $R_{0}(D,0)$ is information theoretically optimal if perception is evaluated in the strong sense of total variation $\delta_{\mathsf{TV}}(P_{X^{n}},P_{\hat{X}^{n}})\to 0$. Under the squared Wasserstein perception (strong-sense), whether $R_{p}(D,0)$ is also the lower bound, or if better rates can be achieved, is an open question. As discussed in Sec. 2.2, this perception notion is more aligned with those used in practice compared to weak-sense or total variation (strong-sense).
@@
PD-LTC does not use a shared dither; we cannot make use of classical results (Zamir and Feder, 1996) to analyze the statistical behavior of latent LQ (as opposed to Thm. 4.4). Instead, Thm. 4.6 relies on lattice Gaussian methods (Ling and Belfiore, 2014). Using scalar transforms, $Q_{\Lambda}(g_{a}(\bm{x}))$ behaves like a lattice Gaussian. The additive private dither $\bm{u}$ (which becomes Gaussian-like) makes the reconstruction approximately Gaussian. It then suffices for $g_{s}$ to scale $\hat{X}^{n}$ to impose the desired variance $\sigma^{2}$.
@@ 5 Experimental Results
We denote $n$ the source dimension and $n_{L}$ the latent space dimension. We denote LTC as NTC when the lattice is chosen to be the integer lattice, i.e., $\Lambda=\mathbb{Z}_{n_{L}}$.
Experimental setup. We train the PD-LTC, SD-LTC, and QSD-LTC models using the objectives in (2), (11), and (15) respectively. For the rates reported at test time, we use the (6) version of $H(\bm{c}|\bm{u})$ for SD-LTC, $\operatorname{\mathbb{E}}_{\bm{y}}[-\log p_{\hat{\bm{y}}}(\hat{\bm{y}})]$ for PD-LTC, and $H(\bm{c}|\hat{\bm{u}})$ for QSD-LTC, all which require hard quantization. These rates are cross-entropy upper bounds on the true entropy, due to the learned $p_{\bm{y}}$ density. We use MSE distortion $\operatorname{\Delta}(\bm{x},\hat{\bm{x}})=\frac{1}{n}\|\bm{x}-\hat{\bm{x}}\|_{2}^{2}$. For perception, to obtain reliable estimates in higher dimensions, we use squared sliced Wasserstein distance (Bonneel et al., 2015) of order 2, $\delta(P_{\bm{x}},P_{\hat{\bm{x}}})=\frac{1}{n}\mathsf{SW}_{2}^{2}(P_{\bm{x}},P_{\hat{\bm{x}}})$. During training, we use the straight-through estimator with hard quantization to generate the reconstructions $\hat{\bm{x}}$.
@@ 5.1 Synthetic Sources
We first evaluate the i.i.d. Gaussian source of dimension $n=8$. We set $n_{L}=8$, $g_{a}$ and $g_{s}$ to be linear functions, as the constructions in Thm. 4.4 and Thm. 4.6 suggest this to be sufficient for optimality, and use the $\mathbb{Z}_{8}$ and $E_{8}$ lattices. To cover the RDP tradeoff, we sweep a variety of $\lambda_{1},\lambda_{2}$ values. In Fig. 4, we plot the equi-perception curves with $P=0$ to compare the methods under the perfect realism setting. As shown, for a fixed amount of shared randomness, a more efficient lattice improves performance. Analogously, for a fixed lattice, more shared randomness improves performance. We additionally evaluate QSD-LTC with finite shared randomness. The fine lattice $\Lambda_{f}$ in Def. 3.4 is set to be self-similar with $\Lambda$, with a nesting ratio of $3$. This results in $R_{c}=\log 3$ bits per dimension. As shown, the QSD-LTC performance is nearly able to achieve that of SD-LTC, despite not having infinite shared randomness; this verifies that QSD-LTC allows one to interpolate between SD-LTC and PD-LTC. At low rates, the LTC with $E_{8}$ has some suboptimality; this is due to the Monte-Carlo estimation when computing latent likelihoods (Lei et al., 2025).
When compared to the fundamental limits described in Sec. 4, we see that the PD models are lower bounded by $R_{p}(D,0)$; the SD models are are lower bounded by the RDP function $R(D,0)$ but outperform $R_{p}(D,0)$ when the rate is not too small. Furthermore, PD-LTC with $E_{8}$ already matches $R_{0}(D,0)$ at lower rates, supporting our theoretical finding that $R_{0}(D,0)$ is not the private randomness fundamental limit under the perception constraint in (24). Although the fundamental RDP limits should be interpreted operationally with perception measured as Wasserstein and not sliced Wasserstein, we empirically verify that on Gaussians, sliced Wasserstein is faithful to Wasserstein in Appendix. D. The performance of $n=1$ RCC (i.e., each dimension of $\bm{x}$ is compressed with RCC separately) and $n=8$ RCC is shown in Fig. 9(a). Performance improves with increasing dimension. However, the 8-dimensional RCC is outperformed by SD-LTC, and additionally has complexity exponential in dimension and rate; SD-LTC does not suffer the same high complexity. We show the RDP achieved by deterministic NTC in Fig. 9(b); at low rates the lack of randomness prevents it from enforcing the perception constraint. At larger rates, its performance coincides with SD-NTC.
Figure 4: Effect of lattice choice and shared randomness on RDP achieved at $P=0$ on the Gaussian source.
@@ 5.2 Real-World Sources
Figure 5: Effect of lattice choice and shared randomness on RDP tradeoff, on Physics.
Figure 6: Effect of lattice choice and shared randomness on RDP tradeoff, on Speech.
Figure 7: Comparing PD-LTC, QSD-LTC-$\Gamma$, and SD-LTC on Physics. $\Gamma$ denotes the nesting ratio.
Figure 8: Comparing PD-LTC, QSD-LTC-$\Gamma$, and SD-LTC on Speech. $\Gamma$ denotes the nesting ratio.
We use the Physics and Speech datasets (Yang and Mandt, 2022), where the former contains physics measurements of dimension $n=16$, and the latter contains audio signals of dimension $n=33$. We use a latent dimension of $n_{L}=8$ and $n_{L}=16$ respectively. We use the integer lattice for NTC models; for Speech, we use the $E_{8}$ lattice, and for Physics, we use the $\Lambda_{16}$ lattice (Barnes and Wall, 1959). We use MLPs for $g_{a}$ and $g_{s}$ of depth 3, hidden dimension 100, and softplus nonlinearities. The corresponding RDP tradeoff achieved on Physics is shown in Fig. 5, and Speech is shown in Fig. 6. Due to lack of randomness, deterministic NTC and LTC are unable to enforce the perception constraint at lower rates, no matter how large $\lambda_{2}$ in (2) is set. This demonstrates the benefits of lattices and shared randomness described in the prior sections translate to real-world sources that require nonlinear transforms. Similar to the Gaussian case, performance improves with better lattices and increased shared randomness. For QSD-LTC, we use self-similar nested lattices with a nesting ratios of 2 and 3, corresponding to $R_{c}=\log 2=1$ and $R_{c}=\log 3\approx 1.58$ respectively. Shown in Figs. 7, 8, we see that performance increases with more shared randomness. A full comparison of all models is shown in Fig. 10 for Physics and Fig. 11 for Speech. Overall, SD-LTC is the best performing variant, achieving the lowest rate-distortion and rate-perception tradeoffs. For a fixed lattice choice, QSD-LTC, with $\log 3$ bits of shared randomness, can nearly achieve the performance of SD-LTC, which uses infinite shared randomness.
@@ 5.3 Ablation Study
We perform an ablation study on the lattice choice and the training of PD-LTC with STE or the noisy proxy. We use the $D_{n}^{*}$ lattice for SD-LTC in Fig. 12 for the Speech dataset. As shown, its performance lies between that of the integer and Barnes-Wall lattices, which aligns with the fact that the $D_{n}$ packing efficiency lies between those two lattices. This supports our result in Thm. 4.4 that performance is optimal when lattices have lower NSM. For the noisy proxy, we use it to train PD-LTC (as is done in the literature), but this may result in a train/test mismatch, since unlike SD-LTC, its rate does not equal the rate under hard quantization. We train PD-LTC with hard quantization, using STE for backpropagation. Shown in Fig. 13, the resulting performance is very similar to that of the noisy proxy.
6 Conclusion
We investigate low-complexity methods that perform well for the RDP tradeoff. We propose to combine dithered lattice quantization with neural compression, which supports different amounts of shared randomness. We theoretically analyze their performance under the infinite and no shared randomness settings, and show that SD-LTC achieves optimality on the Gaussian source. We empirically verify that performance improves with increased shared randomness and improved lattice efficiency. Future work may include expanding the range of the rate of shared randomness in QSD-LTC, as well as its theoretical analysis.

### v2

@@ 4 Achieving the Fundamental Limits
We now theoretically analyze the performance of SD-LTC and PD-LTC on the Gaussian source, and describe the RDP tradeoff asymptotically achievable by SD-LTC and PD-LTC. While the operational rate and distortion are given by the per-dimension versions of those in Sec. 3, the operational perception used is per-dimension squared 2-Wasserstein distance. We leave full proofs to Sec. D.
@@ 4.1 Infinite Shared Randomness
Let $P_{X}=\mathcal{N}(0,\sigma^{2})$, $\operatorname{\Delta}(x,\hat{x})=(x-\hat{x})^{2}$, and $\delta(\mu,\nu)=W_{2}^{2}(\mu,\nu)$ be squared 2-Wasserstein distance. Then
@@
The following theorem shows that $R(D,P)$ is achievable with SD-LTCs, and is an extension of Li et al. (2011), whose authors addressed the $P=0$ case.
Theorem 4.3 (Optimality of SD-LTC for Gaussian source).
Let $X_{1},X_{2},\dots\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\mathcal{N}(0,\sigma^{2})$. For any $P\in[0,\sigma^{2}],D\in[0,2\sigma^{2}]$, there exists a sequence of SD-LTCs $\{(g_{a}^{(n)},g_{s}^{(n)},\Lambda^{(n)})\}_{n=1}^{\infty}$ such that
@@
$\lim_{n\rightarrow\infty}\frac{1}{n}W_{2}^{2}(P_{X^{n}},P_{\hat{X}^{n}})\leq P,$ (13)
@@
The proof of Thm. 4.3 relies on a sphere-bound-achieving sequence of lattices with scalar transforms $g_{a}$, $g_{s}$. As dimension grows, the dither $\bm{u}$ becomes Gaussian-like, and the latent dithered LQ acts like an additive Gaussian channel (Zamir and Feder, 1996), imposing a joint Gaussian relationship between $X$ and $\hat{X}$ as desired by the RDP solution; see Remark 4.2 and Fig. 3.
@@ 4.2 No Shared Randomness
We now consider the case when the encoder and decoder do not have access to any shared randomness. For simplicity and ease of presentation, we consider the regime of near-perfect and perfect perception, corresponding to $P<\epsilon$ for any $\epsilon>0$ and $P=0$, respectively. This is the only regime of perception where lower bounds on the RDP achievable under no shared randomness are known (Saldi et al., 2015; Wagner, 2022; Chen et al., 2022; Yan et al., 2021), which corresponds to $R\left(\nicefrac{{D}}{{2}},\infty\right)$, and evaluates to $\frac{1}{2}\log\frac{2\sigma^{2}}{D}$ on the i.i.d. Gaussian source. The following theorem shows that $R\left(\nicefrac{{D}}{{2}},\infty\right)$ is achievable with PD-LTCs under near-perfect perception, which can be easily extended to perfect perception (see Remark D.8). As discussed in Sec. 2, $R(D,0)<R(\nicefrac{{D}}{{2}},\infty)$; see Fig. 4 for a visualization.
Figure 3: In the latent space, SD-LTC (left) becomes AWGN-like. PD-LTC (right) models the sum of a lattice Gaussian and Gaussian; $s$ must be large enough for the sum to be Gaussian.
Theorem 4.5 (Optimality of PD-LTC for Gaussian source).
@@
$\lim_{n\rightarrow\infty}\frac{1}{n}H(Q_{\Lambda^{(n)}}(g_{a}^{(n)}(X^{n})))\leq R\left(\nicefrac{{D}}{{2}},\infty\right),$ (14)
$\lim_{n\rightarrow\infty}\frac{1}{n}\mathbb{E}[\|X^{n}-\hat{X}^{n}\|^{2}]\leq D,$ (15)
@@
where $\hat{X}^{n}=g_{s}^{(n)}(Q_{\Lambda^{(n)}}(g_{a}^{(n)}(X^{n}))+s\bm{u})$, $\bm{u}\sim\mathrm{Unif}(\mathcal{V}(\Lambda^{(n)}))$, and $s=\frac{\sigma}{\sqrt{\sigma^{2}-\nicefrac{{D}}{{2}}}}$.
@@
PD-LTC does not use a shared dither; we cannot make use of classical results (Zamir and Feder, 1996) to analyze the statistical behavior of latent LQ as an additive channel (as opposed to Thm. 4.3). Instead, Thm. 4.5 relies on lattice Gaussian coding methods (Ling and Belfiore, 2014); see Sec. D.3 for details. Using scalar transforms, $Q_{\Lambda}(g_{a}(\bm{x}))$ behaves like a lattice Gaussian. The additive private dither $s\bm{u}$ (which becomes Gaussian-like) makes the reconstruction Gaussian-like (Fig. 3). It then suffices for $g_{s}$ to scale $\hat{X}^{n}$ to impose the desired variance $\sigma^{2}$. In classical rate-distortion, dithered LQ and lattice Gaussian coding can both achieve optimality, so their differences are primarily practical (e.g., restrictions on lattice families, availability of a shared dither). Thms. 4.3, 4.5 show that with a perception constraint, the two proof techniques lead to different fundamental limits as well.
@@
We note that the choice of $s={\sigma}/{\sqrt{\sigma^{2}-\nicefrac{{D}}{{2}}}}>1$ depends on the regime of the rate-distortion tradeoff; it is approximately $1$ for small $D$, and becomes large for $D\rightarrow 2\sigma^{2}$. At low rates, this means the effective dither $s\bm{u}$ “leaks” outside the lattice cell rather significantly. This is required to ensure the perception term vanishes in (16). Specifically, $s\bm{u}$ needs to be large enough relative to the quantization steps of $Q(g_{a}^{(n)}(X^{n}))$ for $Q_{\Lambda^{(n)}}(g_{a}^{(n)}(X^{n}))+s\bm{u}$ to approximate a Gaussian of covariance $\sigma^{2}I_{n}$ arbitrarily accurately. The nature of this approximation is based on the flatness factor (Ling et al., 2014, Def. 5) used in lattice Gaussian coding, which is elaborated on in Remark D.7.
@@ 5 Experimental Results
Figure 4: Effect of lattice choice and shared randomness on Gaussians at $P=0$.
Experimental setup. We denote $n$ the source dimension and $n_{L}$ the latent space dimension. We denote LTC as NTC when the lattice is chosen to be the integer lattice, i.e., $\Lambda=\mathbb{Z}_{n_{L}}$. We train the PD-, SD- and QSD-LTC models using their RDP objectives discussed in Sec. 3. For the rates reported at test time, we use the (4) version of $H(\bm{c}|\bm{u})$ for SD-LTC, $\operatorname{\mathbb{E}}_{\bm{y}}[-\log p_{\hat{\bm{y}}}(\hat{\bm{y}})]$ for PD-LTC, and $H(\bm{c}|\hat{\bm{u}})$ for QSD-LTC; all require hard quantization. These rates are cross-entropy upper bounds on the true entropy, due to the learned $p_{\bm{y}}$ density. We use MSE distortion $\operatorname{\Delta}(\bm{x},\hat{\bm{x}})=\frac{1}{n}\|\bm{x}-\hat{\bm{x}}\|_{2}^{2}$. For perception, to obtain reliable estimates in higher dimensions, we use squared sliced Wasserstein distance (Bonneel et al., 2015) of order 2, $\delta(P_{\bm{x}},P_{\hat{\bm{x}}})=\frac{1}{n}\mathsf{SW}_{2}^{2}(P_{\bm{x}},P_{\hat{\bm{x}}})$. During training, we use the straight-through estimator with hard quantization. See Sec. B for further details on architecture/training.
@@ 5.1 Synthetic Sources
We first evaluate the i.i.d. Gaussian source of dimension $n=8$. In Fig. 4, we plot the equi-perception curves with $P=0$ to compare the methods under the perfect realism setting. As shown, for a fixed amount of shared randomness, a more efficient lattice improves performance. Analogously, for a fixed lattice, more shared randomness improves performance. We additionally evaluate QSD-LTC with finite shared randomness. The fine lattice $\Lambda_{f}$ in Def. 3.4 is set to be self-similar with $\Lambda$, with a nesting ratio of $3$. This results in $R_{c}=\log 3$ bits per dimension. As shown, the QSD-LTC performance is nearly able to achieve that of SD-LTC, despite not having infinite shared randomness; this verifies that QSD-LTC allows one to interpolate between SD-LTC and PD-LTC. At low rates, the LTC with $E_{8}$ has some suboptimality; this is due to the Monte-Carlo estimation when computing the integral for latent likelihoods (Lei et al., 2025).
When compared to the fundamental limits described in Sec. 4, we see that the PD models are lower bounded by $R(\nicefrac{{D}}{{2}},\infty)$, and the SD models are are lower bounded by the RDP function $R(D,0)$ but outperform $R(\nicefrac{{D}}{{2}},\infty)$ when the rate is not too small. Although the fundamental RDP limits should be interpreted operationally with perception measured as Wasserstein and not sliced Wasserstein, we empirically verify that on Gaussians, sliced Wasserstein is faithful to Wasserstein in Appendix. E. The performance of $n=1$ RCC (i.e., each dimension of $\bm{x}$ is compressed with RCC separately) and $n=8$ RCC is shown in Fig. 7(a). Performance improves with increasing dimension. However, the 8-dimensional RCC is outperformed by SD-LTC, and additionally has complexity exponential in dimension and rate; SD-LTC does not suffer the same high complexity. We show the RDP achieved by deterministic NTC in Fig. 7(b); at low rates the lack of randomness prevents it from enforcing the perception constraint. At larger rates, its performance coincides with SD-NTC.
@@ 5.2 Real-World Sources
Figure 5: LTC, PD-LTC, and SD-LTC on real-world sources; R-D (top) and R-P (bottom).
Figure 6: QSD-LTC-$\Gamma$ ($\Gamma$ the nesting ratio) on real-world sources; R-D (top) and R-P (bottom).
We use MNIST (Lecun et al., 1998), Physics and Speech datasets (Yang and Mandt, 2022). These contain grayscale images of dimension $28\times 28$, physics measurements of dimension $16$, and audio signals of dimension $33$ respectively. We use a latent dimension of $n_{L}=8$ for the first two and $n_{L}=16$ for Speech. We use the integer lattice for NTC models; for MNIST/Physics, we use the $E_{8}$ lattice, and for Speech, we use the $\Lambda_{16}$ lattice (Barnes and Wall, 1959). The corresponding RDP tradeoff is shown in Fig. 5. Due to lack of randomness, deterministic NTC and LTC are unable to enforce the perception constraint at lower rates, no matter how large $\lambda_{2}$ in (2) is set. Similar to the Gaussian case, performance improves with better lattices and increased shared randomness, demonstrating the benefits of lattices and shared randomness described in the prior sections translate to real-world sources that require nonlinear transforms. For QSD-LTC, we use self-similar nested lattices with a nesting ratios of 2 and 3, corresponding to $R_{c}=\log 2=1$ and $R_{c}=\log 3\approx 1.58$ respectively. Performance increases with $R_{c}$ (Fig. 6). A full comparison across all models/datasets is shown in Figs. 8, 9 and 10. Overall, SD-LTC achieves the RDP tradeoffs. For a fixed lattice, QSD-LTC, with $R_{c}=\log 3$, can nearly achieve the performance of SD-LTC, which uses $R_{c}=\infty$.
@@ 5.3 Ablation Study
We perform an ablation study on lattice choice and training of PD-LTC with STE or the noisy proxy. We use the $D_{n}^{*}$ lattice for SD-LTC in Fig. 12 on Speech. Its performance lies between integer and Barnes-Wall lattices, which aligns with the fact that the $D_{n}$ packing efficiency lies between those two lattices. This supports our result in Thm. 4.3 that performance is optimal when the lattices pack space more efficiently. For the noisy proxy, we use it to train PD-LTC, but this may result in a train/test mismatch, since unlike SD-LTC, the noisy proxy does not equal the rate under hard quantization. A comparison of the two in Fig. 12 shows there is not much difference in the resulting performance.
6 Conclusion and Limitations
We investigate low-complexity, high-performing compressors for the RDP tradeoff. We propose combining dithered LQ with neural compression, which supports different amounts of shared randomness. Under infinite and no shared randomness, we show SD-LTC and PD-LTC achieve optimality on the Gaussian source. We empirically verify that performance improves with increased shared randomness and improved lattice efficiency across synthetic and real-world data. Future work may address: expanding the range of $R_{c}$ in QSD-LTC, its theoretical analysis, and extension to SOTA image compression architectures, which would require careful integration of random dithering with hyperprior models that implicitly use a deterministic dither.